A body `A` of mass 4 on collision exerts a force on another body B on mass 10. If the acceleration produced in B is 10ms^-2,then what`s the acceleration of A?

Grade:10

1 Answers

Arun
6 years ago
According to Newton's third law of motion
Action = -(reaction)
Hence
F1 = - F2
m1* a1 = - m2 *a2
a1 = - 10* 10/4
a1 = -25 m/sec²
Magnitude wise acceleration in A = 25 m/sec²